• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Nilanga

    Log entry time 23:55:35 on February 20,1999

    Entry number 14999

    This entry is a followup to: 14861

    keyword=Sintillator efficiency

    The efficiency of the electron spectrometer S2 left phototube # 3 was found to be low this morning. The HV was raised to 2180 V. Bogdan pointed out that the high voltage limit for these tubes are around 2300 V. So this tube ( and several other tubes) may need to be replaced soon.